OpenStudy (anonymous):

Simplify \(\frac{cos^{2}a - cot^{2} a+1}{sin^{2}a+tan^{2}a -1}\)

OpenStudy (solomonzelman):

I would start from simplifying the denominator \(\Large\color{orangered}{ sin^2a +tan^2a+1 }\) \(\Large\color{orangered}{ sin^2a +\frac{sin^2a}{cos^2a}+1 }\) \(\Large\color{orangered}{ \frac{sin^2a~cos^2a}{cos^2a} +\frac{sin^2a}{cos^2a}+\frac{cos^2a}{cos^2a} }\) \(\Large\color{orangered}{ \frac{sin^2a~cos^2a+ sin^2a+ cos^2a}{cos^2a} }\) then the numerator, \(\Large\color{orangered}{ cos^2a +cot^2a+1 }\) \(\Large\color{orangered}{ cos^2a +\frac{cos^2a}{sin^2a}+1 }\) \(\Large\color{orangered}{ \frac{cos^2a~sin^2a}{sin^2a} +\frac{cos^2a}{sin^2a}+\frac{sin^2a}{sin^2a} }\) \(\Large\color{orangered}{ \frac{sin^2a~cos^2a+ sin^2a+ cos^2a}{cos^2a} }\)

OpenStudy (solomonzelman):

yeah, the denominator and numerator are equal. I switched the order a little but a+b=b+a so it doesn't matter. The answer is 1.

OpenStudy (anonymous):

Thank you @SolomonZelman :D

OpenStudy (solomonzelman):

Anytime